OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)). answer choices -9x - 2 -9x + 5 -9x + 25 -9x + 27

OpenStudy (zzr0ck3r):

\(f(\color{blue}{g(x)}) = f(\color{blue}{-x+3})=9(\color{blue}{-x+3})-2=-9x+27-2=-9x+25\)

OpenStudy (zzr0ck3r):

all I did was plug in the function \(g(x)\) where ever I saw a \(x\) in \(f(x)\)

OpenStudy (anonymous):

but wouldn't that change the function

OpenStudy (zzr0ck3r):

we are taking two functions and making a third function out of them

OpenStudy (zzr0ck3r):

we have three operations with functions \(f+g. f*g, f\cdot g\) the last being composition

OpenStudy (zzr0ck3r):

so yes just as adding two functions changes things, so does composing two function.

OpenStudy (zzr0ck3r):

just like 3+4=7 changes 3, and 4 into 7

OpenStudy (anonymous):

ohh i get it now thank you!!

OpenStudy (zzr0ck3r):

have you seen the human centipede movie? or know about it? (if not do NOT google it)

OpenStudy (zzr0ck3r):

anyway this is how I explain composition of functions to students:)